Write the expression for the perimeter of a rectangle with a length that is 4 centimeters longer than three times its width

Answer by clearblueskai(22)

Perimeter: P = 2(l+w)
Length is 4 centimeters longer than thrice of its width:
l= 4+3w
P= 2((4+3w)+w)
P= 8+8w